A tablet weighing 0.6525 g contains approximately 300 mg paracetamol. Calculate 

the mass of tablet that must be dissolved in water and diluted to 500.0 mL to prepare 

a solution containing 25 ppm (m/v) paracetamol


1
Expert's answer
2021-06-28T08:15:23-0400

0.6525×1000 = 652.5mg

652.5 - 300 = 352.5mg

=300/352 × 500

= 426.14mg